OpenStudy (anonymous):

Solve ln (x-4) < 6 and solve the inequality 3^2n less then or equal to 1/81

OpenStudy (anonymous):

\[\ln(x)=y\iff e^y=x\] and since both are strictly increasing \[\ln(x)<y\iff x<e^y\]

OpenStudy (anonymous):

in your example \[\ln(x-4)<6\iff x-4<e^6\]

OpenStudy (anonymous):

what i put down as my answer before was x < e^6 +4 is that wrong?

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

so that answers correct? Thanks!

OpenStudy (anonymous):

yes as for the next one, note that \(\frac{1}{81}=3^{-4}\)

OpenStudy (anonymous):

so is that the answer or how i would begin to solve it?

OpenStudy (anonymous):

\[3^{2n}<3^{-4}\iff 2n<-4\iff n<-2\]

OpenStudy (anonymous):

I'm a bit confused, so the answer would be n <-2 ?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yes, that is the answer

OpenStudy (anonymous):

Thanks for all of your help!